Circular No. 7591

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001X IN NGC 5921 W. Li, University of California at Berkeley, communicates that Y. Fan,

Y. L. Qiu, and J. Y. Hu, on behalf of Beijing Astronomical Observatory (BAO) Supernova Survey, report the discovery of an apparent supernova (mag ∼ 17.0) found on an unfiltered image taken with the BAO 0.6-m telescope on Feb. 27.8 UT. SN 2001X is located at α = 15^{h}21^{m}55^{s}.45, δ = +5^{o}03′42′′.1 (equinox 2000.0), which is 15′′.5 west and 32′′.4 south of the nucleus of NGC 5921. SN 2001X was confirmed on an unfiltered image taken by M. Schwartz on Mar. 2.5 with the Tenagra Observatory 0.5-m automatic telescope, from which Li measured the position end figures to be 55^{s}.46 ± 0^{s}.03, 43′′.1 ± 0′′.2, and the magnitude of the supernova to be 15.2. Images of the same field taken with the BAO 0.6-m telescope on Feb. 6.9 (limiting mag ∼ 17.5) and with the Tenagra 0.5-m telescope on Feb. 17.5 (limiting mag ∼ 18.0) showed nothing at this position. Li remarks that a normal type-Ia supernova in this galaxy might reach mag ∼ 13.

V4643 SAGITTARII = NOVA SAGITTARII 2001 N. N. Samus, Institute of Astronomy, Moscow, informs us that the

designation V4643 Sgr has been given to this nova (cf. IAUC 7589, 7590).

The USNO-A2.0 catalogue contains a red star (also visible on the Digital Sky Survey) within 1′′ of the nova’s position, at position end figures 40^{s}.41, 15′′.4 (cf. IAUC 7589), with the blue and red magnitudes listed as 17.4 and 15.8, respectively. Additional visual magnitude estimates: Mar. 1.249 UT, 10.2 (J. G. de S. Aguiar, Campinas, Brazil); 1.642, 10.4 (J. Bedient, Honolulu, HI). T. Kato, Kyoto University, reports that Y. Nakamura (Suzuka, Mie) found no sign of the nova to limiting mag ∼ 11.1 on a Tri-X exposure taken on Feb. 20.852 with a 200-mm f/4 lens.

SUPERNOVA 2001C IN PGC 19975 CCD photometry, unfiltered unless otherwise noted: Jan. 12.783 UT,

R = 14.96 (K. Hornoch, Lelekovice, Czech Republic); 13.824, R = 15.00 (Hornoch); 13.89, V = 15.1 (D. Hanžl, Brno, Czech Republic); 14.822, R = 15.00 (Hornoch); 28.826, R = 15.12 (Hornoch); 31.767, 15.04 (M. Martignoni, Busto Arsizio, Italy); Feb. 2.743, 14.95 (Martignoni); 2.805, R = 15.30 (Hornoch); 11.853, R = 15.76 (Hornoch); 13.800, R = 15.86 (Hornoch).

2001 March 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green